The angle of elevation from a point on the ground to the top of a pyramid is 3434degrees°4040​' The angle of elevation from a point 134134 feet farther back to the top of the pyramid is 2323degrees°4040​'. Find the height of the pyramid.
Answer by Boreal(15235)

The tangent is the height/distance
tan 34d40'=height/D
height=D*0.6916
--------------------
tangent further back is height/(D+134)
tangent 23d40'=height/D
height= (D+134)*0.4383
-------------------
Those two are equal
0.6916D=0.4383D+58.73, the product of 134*0.4383
0.2533D=58.73
D=231.86 feet
Now can get the height, which is D*0.6916 or 160.4 feet ANSWER
Check with second tangent 23d40m=160.4/365.86=23.67d, checks.